OpenStudy (anonymous):

Two lines through the point (-1,3) are tangent to the curve x^2+4y^2-4x-8y+3=0. find the equation.

OpenStudy (anonymous):

I have been workin on this last night but can't get the right slope

OpenStudy (anonymous):

the correct slope is suppose to be -11/4 and -11/44 but can't get it to be like that. I used implicit differentiation.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Hero @markaskingalexandria1 @m3m3man1999 @MahoneyBear @MathHater82 @m&msdoodle @marigirl @e.s.b @Empty @e.v.

OpenStudy (anonymous):

pls don't tag me ;-;

OpenStudy (anonymous):

here is what i have done so far but the slopes i got are -11/2 etc...

OpenStudy (anonymous):

@m&msdoodle @Hero @theopenstudyowl @pebbles_xx3 @PlasmaFuzer @plohrb @plzzhelpme @pooja195 @proud_yemeniah_ @pacely.chitlen.98

hero (hero):

Calculus is not necessarily needed for this.

OpenStudy (anonymous):

but the only point given is (-1,3)

OpenStudy (anonymous):

the equation given is a conic and it is an ellipse

OpenStudy (anonymous):

how would you solve it sir?

hero (hero):

I take that back. At first I thought it was a circle.

OpenStudy (anonymous):

any hints to get the slope? which is the dy/dx?

OpenStudy (marigirl):

@jim_thompson5910

OpenStudy (marigirl):

@misty1212

OpenStudy (marigirl):

would you happen to have the original question? could u post it?

OpenStudy (anonymous):

Two lines through the point (-1,3) are tangent to the curve x^2+4y^2-4x-8y+3=0. find the equation.

jimthompson5910 (jim_thompson5910):

@RAY_OMEGA you should get \[\Large \frac{dy}{dx} = \frac{-2x+4}{8y-8} = \frac{-x+2}{4y-4}\]

OpenStudy (anonymous):

Yeah my dy/dx is x-2/4(y-1)

jimthompson5910 (jim_thompson5910):

Let point P be the point (-1,3) Let Q be some point (x,y) on the elliptical curve Find the slope of line PQ \[\Large m = \frac{y_2-y_1}{x_2-x_1}\] \[\Large m = \frac{y-3}{x-(-1)}\] \[\Large m = \frac{y-3}{x+1}\] ------------------------------------------------------- set that slope equal to dy/dx \[\Large m = \frac{y-3}{x+1}\] \[\Large \frac{dy}{dx} = \frac{y-3}{x+1}\] \[\Large \frac{-x+2}{4(y-1)} = \frac{y-3}{x+1}\] at this point I'm stuck. I have a feeling you have to solve `x^2+4y^2-4x-8y+3=0` for y, then replace each y with that expression in terms of x, but that looks like it will get very messy real fast.

OpenStudy (anonymous):

Sorry for no reply I had to attend class

OpenStudy (anonymous):

So here is what I did

OpenStudy (anonymous):

After I got the dy/dx of the given curve I used at point slope formula plugging in (-1,3) as the points x and y and dy/dx asy slope

OpenStudy (anonymous):

Hey guys i already solved the problem here just gonna leave this open if you want to know the correct answer